OpenStudy (anonymous):

the equation 9x+2y=8 describes a line and is written in standard form. a) solve for variable y b) use this result to find y with x=-2

OpenStudy (anonymous):

9x + 2y = 8 -9x 2y = 8 -9x -- 2 y = 8 - 9x ------ 2 y = 4 -9/2x

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

part of it yes. do you understand what i did to get y?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

great, now for the second part, all you have to do is put -2 where x is.. so.. y = 4 - 9/2(-2)

OpenStudy (anonymous):

ok, finish

OpenStudy (anonymous):

why don't you try it and tell me what you get and i will tell you if we got the same answer or not

OpenStudy (anonymous):

yes i got the answer

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

y=4-9/2(-2)

OpenStudy (anonymous):

... you have to actually do it. multipy, subtract, add, etc. until there is only one number that is equal to y

OpenStudy (anonymous):

i'm blind

OpenStudy (anonymous):

excuse me? are you speaking metaphorically?

OpenStudy (anonymous):

i can type normal